OpenStudy (anonymous):

find the eqaution of the tangent plane at the point (2,2) on the surface z = sin (x^y)

OpenStudy (anonymous):

1)first find the derivative of the following equation = m =equation of the slope 2) substitute x to find the slope 3) substitute the rest of your given in the following equation to find the equation of the tangent line : yp - y = m(x - xp) and simply solve ^_^ clearer now?

OpenStudy (anonymous):

the question is not asking about the tangent line in 2D, but about the tangent plane on a surface

OpenStudy (anonymous):

is it \[z=\sin (xy)\]

OpenStudy (anonymous):

lol sorry ^^"

OpenStudy (anonymous):

Would this involve the del operator (sum of partial derivatives)?

OpenStudy (anonymous):

let \[f(x,y,z)=z-\sin(xy)\] grad f(x,y,z) = \[<-ycos(xy),-xcos(xy),1>\] grad f(2,2,z)= <-2cos(4),-2cos(4),1>

OpenStudy (anonymous):

the equation of the tangent plane will be: -2cos(4)(x-2)-2cos(4)(y-2)+(z-sin(4))=0 simplify!!

OpenStudy (anonymous):

I gotta go now